# Uplink-Downlink Duality for Beamforming in Integrated Sensing and Communications

### v1

@@ IV-B Uplink-Downlink Duality
Here, $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0$ in constraint (57d) means that $\mathbf{M}_{\mathbf{U}}$ is invertible and $\mathbf{M}_{\mathbf{U}}^{-1}$ has nonnegative entries.
@@
Figure 3: Uplink-downlink duality holds for downlink weighted power minimizstion with non-PSD $(\boldsymbol{\lambda}\mathbf{I}-\mathbf{Q}_{\beta})$, if $(\lambda,\boldsymbol{\beta})$ is admissible. The normalized optimal beamformers $\mathbf{v}_{k}$ for the downlink (left) are the same as the optimal beamformers $\mathbf{u}_{k}$ for the uplink (right) with an $M$-matrix constraint.
@@
$\mathrm{s.t.}$~~~~ $\mathbf{M}_{\mathbf{U}}\mathbf{p}\geq\sigma^{2}\mathbf{1},$ (62b)
@@
The matrix $\mathbf{M}_{\mathbf{U}}$ has the specific form of having positive diagonal elements and having off-diagonal elements that are less than or equal to zero. We claim that as long as $\mathbf{U}$, along with some $\mathbf{p}$, is a feasible solution to (62), the matrix $\mathbf{M}_{\mathbf{U}}$ must belong to a class of so-called M-matrices [51]. There are many equivalent definitions of the $M$-matrix; several of these are listed in Appendix C. The reason that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix is because the feasibility of (62b)-(62c) implies the existence of a $\mathbf{p}\geq 0$ for which $\mathbf{M}_{\mathbf{U}}\mathbf{p}>0$. This is one of the equivalent definitions of an $M$-matrix.
@@
$\mathbf{p}=\sigma^{2}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{1},$ (63)
@@
We further claim that for any feasible $\mathbf{U}$, the above solution for $\mathbf{p}$ is actually the optimal solution to the inner minimization problem of (62). This would have been a straightforward statement if $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is PSD, in which case the objective of weighted power minimization (62a) would have nonnegative weights, i.e., ${\boldsymbol{\omega}}_{\mathbf{U}}]_{k}\geq 0$, $\forall k$, due to (61). But for this downlink ISAC problem, the weights $\boldsymbol{\omega}_{\mathbf{U}}$ are not necessarily nonnegative. This is because $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is not necessarily PSD; furthermore, the admissibility condition (54) only ensures that the sum power (or equivalently $\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{p}$) is nonnegative, but allows the individual weights to potentially be negative. This makes sense for sensing problems, because the negative weights indicate the benefit of increasing power toward the target directions.
Nevertheless, we can use the condition $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$ to show that for any feasible $\mathbf{U}$, the inner minimization in (62) would always have the optimal solution achieving equality in (62b), and thus the optimal $\mathbf{p}$ is given by (63). To see this, we make a change of variable and define
@@
$\mathrm{subject\;to}$~~~ $\mathbf{\tilde{p}}\geq\sigma^{2}\mathbf{1},$ (65b)
@@
Now, let $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. For any feasible $\mathbf{U}$ (or equivalently $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix), we have $\boldsymbol{\phi}_{\mathbf{U}}\geq 0$ by Lemma 2. Then, we argue that there must be one optimal solution of (65) that achieves equality in (65b). This is because if the inequality is strict for some $\mathbf{\tilde{p}}_{i}$, one can always decrease $\mathbf{\tilde{p}}_{i}$ without violating (65c) (since $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0$), while improving the objective (65a) or at least keeping it the same (since $\boldsymbol{\phi}^{\mathsf{T}}_{\mathbf{U}}\geq 0$). This shows that $\mathbf{\tilde{p}}=\sigma^{2}\mathbf{1}$ is an optimal solution, so an optimal $\mathbf{p}$ is given by (63).
@@
$\underset{\mathbf{U}}{\mathrm{minimize}}$~~~~ $\sigma^{2}\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{1}$ (67a)
@@
where (67b) states that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix. This is needed to ensure that $\mathbf{U}$ is such that a feasible $\mathbf{p}$ exists.
@@
$\underset{\mathbf{U}}{\mathrm{minimize}}\quad\underset{\mathbf{q}}{\min}$~~~~ $\sigma^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (68a)
@@
For the uplink problem, we choose to explicitly include the requirement that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix as in constraint (68d) to match the downlink formulation. Note that this extra constraint would not have been needed if the uplink noise covariance is PSD, because in that case we have $\boldsymbol{\omega}_{\mathbf{U}}\geq 0$, due to definition (61), and the existence of a feasible power allocation $\mathbf{q}$ implies that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix, due to (68b)-(68c).
@@
We now argue that the solution to the inner minimization problem above is again achieved with equality, just as in the downlink. This is because (68b) and (68d) imply that $\mathbf{q}\geq\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$. By Lemma 2, since we assume admissibility, we have that $\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$ must be nonnegative. Now, since the weights in the objective are all positive (i.e., the all-one vector), using the same argument as in the downlink, we conclude that the optimal solution to the inner minimization problem must be
@@
Hence, the uplink problem can now be reduced to
$\underset{\mathbf{U}}{\mathrm{minimize}}$~~~~ $\sigma^{2}\mathbf{1}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$ (70a)
@@
Theorem 2 states that whenever the pair $\left(\lambda,\boldsymbol{\beta}\right)$ is admissible, the downlink solution of problem (52) can be recovered by solving the dual uplink problem (57) with $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ as the noise covariance matrix. More specifically, the optimal combiners for the uplink problem (57) are also the optimal beamforming directions for the downlink problem (52) and vice versa. Figure 3 illustrates the duality between the downlink and uplink systems that give rise to problem (52) and problem (57).
The key difference between this new uplink-downlink duality relationship and the classical duality theory is that the uplink noise covariance $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ here is not necessarily PSD. Thus, there are potentially choices of $\mathbf{u}_{k}$ for which the uplink noise powers are negative. Theorem 2 states that when the additional $M$-matrix constraint is imposed on the uplink beamformers, the uplink-downlink duality continues to hold despite the potentially negative uplink noise powers. This $M$-matrix constraint would not have been needed to establish duality if the uplink noise covariance were PSD.
@@ IV-C AoA Estimation Example
$\mathrm{s.t.}$ $~~\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (71b)
@@
In Figure 4, we show a cut in the admissible set obtained at fixed $\Im\{b\}=1$ across different values of $\lambda$ and $\Re\{b\}$ for a random realization of the channels, where $\alpha\sim\mathcal{CN}(1,1)$ for the sensing channel, and $h_{ij}\sim\mathcal{CN}(0,1)$ for the communication channel. In the figure, the admissible set $\mathcal{A}$ is given by the shaded region. This is to be contrasted with the set of $(\lambda,b)$ values for which $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is PSD, as shown by the dashed curve (which is plotted by setting $\lambda$ to be the largest eigenvalue of $\mathbf{Q}_{b}$).
Theorem 2 states that for any $(\lambda,b)$ outside the admissible set $\mathcal{A}$, the inner minimization in (71) is unbounded. Moreover, for any $(\lambda,b)$ in the admissible region, the inner problem admits an equivalent formulation as an uplink problem with $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ acting as the uplink noise covariance matrix. As seen from the figure, the admissible set $\mathcal{A}$ is strictly larger than the set of $(\lambda,b)$ for which $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is PSD. This underscores the need to develop the new uplink-downlink duality theory for the case where $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is admissible but not PSD.
If we assume that $b$ is fixed and consider only $\lambda$ for the outer optimization, the admissibility set above reduces to a ray, as in [33] for the energy harvesting scenario. Thus, the above formulation encompasses [33] as a special case [^3] However, the analysis in [33] appears to have overlooked the $M$-matrix constraint for the uplink beamformers and the ensuing complication associated with negative noise power in the duality formulation..

### v2

@@ IV-B Uplink-Downlink Duality
and the notation $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0$ in constraint (58d) means that $\mathbf{M}_{\mathbf{U}}$ is invertible and $\mathbf{M}_{\mathbf{U}}^{-1}$ has nonnegative entries.
Finally, the optimal values of the downlink problem (55) and the uplink problem (58) are the same.
Remarks: The above theorem states that whenever the pair $\left(\lambda,\boldsymbol{\beta}\right)$ is admissible, the downlink solution of problem (52) can be recovered by solving problem (58). Before proceeding to the proof, we first justify why problem (58) can be referred to as an uplink problem, despite that $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}$ is not necessarily PSD. The key reason is that an optimal solution to (58) is always such that
$\sum_{i\neq k}q^{*}_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}^{*}_{k}\right|^{2}+{\mathbf{u}^{*}_{k}}^{\textsf{H}}(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\mathbf{u}^{*}_{k}\geq 0,\quad\forall k.$ (60)
In other words, the right-hand side of constraint (58b) is always nonnegative at the optimum. This quantity can be thought of as the combined interference and noise power, hence justifying (58) as an SINR constraint for the uplink.
Note that (60) is trivial to prove if $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succ 0$, but for the ISAC problem, it is not necessarily true that $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succ 0$. Below, we show that whenever $(\lambda,\boldsymbol{\beta})$ is admissible, then (60) always holds.
The first step is to note that (58b) is always satisfied with equality at the optimum, because if this were not true, then we can always decrease some $q_{k}$ without violating the constraint, while improving the objective. This means that
$\frac{q^{*}_{k}}{\gamma_{k}}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{u}^{*}_{k}\right|^{2}=\sum_{i\neq k}q^{*}_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}^{*}_{k}\right|^{2}+{\mathbf{u}^{*}_{k}}^{\textsf{H}}(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\mathbf{u}^{*}_{k},~~\forall k.$ (61)
Now, using the fact that the optimal $\mathbf{q}^{*}$ must satisfy $\mathbf{q}^{*}\geq\mathbf{0}$ as in (58c), we see that the left-hand side of (61) is nonnegative, which implies that the right-hand side is also nonnegative. In other words, (60) is always satisfied at the optimum as claimed earlier.
We use (60) to show that problem (58) can be viewed as an uplink problem. Suppose that (60) is satisfied with strict inequality for all $k$, i.e., $\sum_{i\neq k}q^{*}_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{k}^{*}\right|^{2}+{\mathbf{u}^{*}_{k}}^{\textsf{H}}(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\mathbf{u}^{*}_{k}>0,\forall k.$ Then, constraint (58b) can be rearranged as
$\underset{\mathbf{U},\;\mathbf{q}}{\mathrm{minimize}}$~~ $\sigma_{\text{c}}^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (62a)
$\mathrm{subject\ to}$~ $\frac{q_{k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}}{\sum_{i\neq k}q_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}+\mathbf{u}_{k}^{\textsf{H}}(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\mathbf{u}_{k}}\geq\gamma_{k},\forall k$ (62b)
[…]
It is now readily seen that the left-hand of (62b) is identical to the usual uplink SINR expression with uplink powers $\mathbf{q}_{1},\ldots,\mathbf{q}_{K}$, uplink beamformers $\mathbf{u}_{1},\ldots,\mathbf{u}_{K}$, and noise covariance $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}$. It should be noted that $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}$ may not be a “noise covariance” in a standard sense, as it is not necessarily PSD as discussed earlier. Despite this, we loosely use the term to refer to such a matrix.
Based on the above interpretation of constraint (62b), we can view problem (62) (and hence the equivalent problem (58)) as an uplink beamforming and power control problem, similar to that considered in the classical work [9]. It should be noted, however, that here, there is an additional constraint (58d), This constraint is discussed in more detail later in the proof of Theorem 2.
When the left-hand side of (60) is zero for some $k$, we can still view (58) as an uplink problem. However, if expressed as in (62), the SINR expression for the $k$-th user would have zero in the denominator. In fact, in this case the optimal uplink power $q^{*}_{k}$ is also zero, so we have zero divided by zero for the SINR expression. This peculiar situation can indeed arise, which is why we prefer to work with problem (58) as opposed to (62), as the former is always well-defined.
Based on the preceding discussion, the statement of Theorem 2 is that, when $(\lambda,\boldsymbol{\beta})$ is admissible, the optimal beamformers for the uplink problem (58) are also optimal as downlink beamforming directions for the downlink problem (55) and vice versa. Furthermore, the two problems have the same optimal value. Figure 3 illustrates the duality between the downlink and uplink systems that gives rise to (55) and (58). We now proceed to prove the theorem.
@@
Figure 3: Uplink-downlink duality holds for downlink weighted power minimization with non-PSD $(\boldsymbol{\lambda}\mathbf{I}-\mathbf{Q}_{\beta})$, if $(\lambda,\boldsymbol{\beta})$ is admissible. The normalized optimal beamformers $\mathbf{v}_{k}$ for the downlink (left) are the same as the optimal beamformers $\mathbf{u}_{k}$ for the uplink (right) with an $M$-matrix constraint.
@@
$\mathrm{s.t.}$~~~~ $\mathbf{M}_{\mathbf{U}}\mathbf{p}\geq\sigma_{\text{c}}^{2}\mathbf{1},$ (65b)
@@
The matrix $\mathbf{M}_{\mathbf{U}}$ has the specific form of having positive diagonal elements and having off-diagonal elements that are less than or equal to zero. We claim that as long as $\mathbf{U}$, along with some $\mathbf{p}$, is a feasible solution to (65), the matrix $\mathbf{M}_{\mathbf{U}}$ must belong to a class of so-called M-matrices [52]. There are many equivalent definitions of the $M$-matrix; several of these are listed in Appendix C. The reason that $\mathbf{M}_{\mathbf{U}}$ must be an $M$-matrix is because the feasibility of (65b)-(65c) implies the existence of a $\mathbf{p}\geq 0$ for which $\mathbf{M}_{\mathbf{U}}\mathbf{p}>0$. This is one of the equivalent definitions of $M$-matrix.
@@
$\mathbf{p}=\sigma_{\text{c}}^{2}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{1},$ (66)
@@
We further claim that for any feasible $\mathbf{U}$, the above solution for $\mathbf{p}$ is actually an optimal solution to the inner minimization problem of (65). This would have been a straightforward statement if $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is PSD, in which case the objective of weighted power minimization (65a) would have nonnegative weights, i.e., ${\boldsymbol{\omega}}_{\mathbf{U}}]_{k}\geq 0$, $\forall k$, due to (64). But for this downlink ISAC problem, the weights $\boldsymbol{\omega}_{\mathbf{U}}$ are not necessarily nonnegative. This is because $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is not necessarily PSD. Furthermore, the admissibility condition (54) only ensures that the sum power (or equivalently $\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{p}$) is nonnegative, but allows the individual weights to potentially be negative. This makes sense for sensing problems, because the negative weights indicate the benefit of increasing power toward the target directions.
Nevertheless, we can use the condition $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$ to show that for any feasible $\mathbf{U}$, the inner minimization in (65) would always have an optimal solution achieving equality in (65b), and thus an optimal $\mathbf{p}$ is given by (66). To see this, we make a change of variable and define
@@
$\mathrm{subject\;to}$~~~ $\mathbf{\tilde{p}}\geq\sigma_{\text{c}}^{2}\mathbf{1},$ (68b)
@@
Now, let $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. For any feasible $\mathbf{U}$ (or equivalently $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix), we have $\boldsymbol{\phi}_{\mathbf{U}}\geq 0$ by Lemma 2. Then, we argue that there must be one optimal solution of (68) that achieves equality in (68b). This is because if the inequality is strict for some $\mathbf{\tilde{p}}_{i}$, one can always decrease $\mathbf{\tilde{p}}_{i}$ without violating (68c) (since $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0$), while improving the objective (68a) or at least keeping it the same (since $\boldsymbol{\phi}^{\mathsf{T}}_{\mathbf{U}}\geq 0$). This shows that $\mathbf{\tilde{p}}=\sigma_{\text{c}}^{2}\mathbf{1}$ is an optimal solution, so an optimal $\mathbf{p}$ is given by (66). Note that in the case where $\boldsymbol{\phi}_{\mathbf{U}}$ has zero entries, multiple optimal solutions for $\mathbf{p}$ exist. They all give the same optimal value. However, one of them is always given by (66).
@@
$\underset{\mathbf{U}}{\mathrm{minimize}}$~~~~ $\sigma_{\text{c}}^{2}\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{1}$ (70a)
@@
where (70b) states that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix. This is needed because otherwise a feasible $\mathbf{p}$ would not exist, so such $\mathbf{U}$ cannot be the optimal solution of the downlink.
@@
$\underset{\mathbf{U}}{\mathrm{minimize}}\quad\underset{\mathbf{q}}{\min}$~~~~ $\sigma_{\text{c}}^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (71a)
@@
For the uplink problem, we choose to explicitly include the requirement that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix as in constraint (71d) to match the downlink formulation. Note that this extra constraint would not have been needed if the uplink noise covariance $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\succ 0$, because in that case we have $\boldsymbol{\omega}_{\mathbf{U}}>0$, due to definition (64), and the existence of a feasible power allocation $\mathbf{q}$ satisfying (71b)-(71c) implies that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix.
@@
We now argue that the solution to the inner minimization problem above is achieved when the uplink SINRs (i.e., (58b) or (71b)) are met with equality, just as in the downlink. This is because for every $\mathbf{U}$ and $\mathbf{q}$ satisfying (71b) and (71d) we must have $\mathbf{q}\geq\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$. By Lemma 2, since we assume admissibility, we have that $\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$ must be nonnegative. Now, since the weights in the objective are all positive (i.e., the all-one vector), we conclude that the optimal solution to the inner minimization problem must be
@@
As an side note, the above solution for $\mathbf{q}$ implies that (61) holds, because for $\mathbf{U}=\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}$, (72) becomes $\mathbf{M}^{\mathsf{T}}_{\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}}\mathbf{q}^{*}=\boldsymbol{\omega}_{\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}}$, which is exactly (61).
Plugging the structure (72) into (71), the uplink problem can now be reduced to
$\underset{\mathbf{U}}{\mathrm{minimize}}$~~~~ $\sigma_{\text{c}}^{2}\mathbf{1}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$ (73a)
@@
The key difference between this new uplink-downlink duality relationship and the classical duality theory is that the uplink noise covariance $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ here is not necessarily PSD. Thus, there are potentially choices of $\mathbf{u}_{k}$ for which the uplink noise powers are negative. Theorem 2 states that when the additional $M$-matrix constraint is imposed on the uplink beamformers, the uplink-downlink duality continues to hold despite the potentially negative uplink noise powers.
@@ IV-C AoA Estimation Example
$\mathrm{s.t.}$ $~~\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k$ (74b)
@@
In Figure 4, we show a cut in the admissible set obtained at fixed $\Im\{b\}=1$ across different values of $\lambda$ and $\Re\{b\}$ for a random realization of the channels, where $\alpha\sim\mathcal{CN}(1,1)$ for the sensing channel, and $h_{ij}\sim\mathcal{CN}(0,1)$ for the communication channel. In the figure, the admissible set $\mathcal{A}$ is given by the shaded region. This is to be contrasted with the set of $(\lambda,b)$ values for which $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is PSD, given by the set of $(\lambda,b)$ values to the right of the dashed curve obtained by setting $\lambda$ to be the largest eigenvalue of $\mathbf{Q}_{b}$.
Theorem 2 states that for any $(\lambda,b)$ outside the admissible set $\mathcal{A}$, the inner minimization in (74) is unbounded. Moreover, for any $(\lambda,b)$ in the admissible region, the inner problem admits an equivalent formulation as an uplink problem with $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ acting as the uplink noise covariance matrix. As seen from the figure, the admissible set $\mathcal{A}$ is strictly larger than the set of $(\lambda,b)$ for which $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is PSD. This underscores the need to develop the new uplink-downlink duality theory for the case where $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is admissible but not PSD [^3] If we assume that $b$ is fixed and consider only $\lambda$ for the outer optimization, the admissibility set reduces to a ray, as in [33] for the energy harvesting scenario. Thus, the formulation in this paper encompasses [33] as a special case. However, the analysis in [33] appears to have overlooked the $M$-matrix constraint for the uplink beamformers and the ensuing complication associated with negative noise power in the duality formulation..